OpenStudy (anonymous):

Lincoln and Gemma are looking at the equation ___. Lincoln says that the solution is extraneous. Gemma says the solution is non-extraneous. Is Lincoln correct? Is Gemma correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (anonymous):

\[\sqrt{2x - 3 } = \sqrt{x}\]